[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Jul 12, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Jul 12 20:19:03 UTC 2017


HSU Shift Log for Wed Jul 12 2017 16:18:38 GMT-0400 (EDT)

Shift: H

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: Both for the 15Z and 21Z advisory, it took 3 tries sending the  
compute data for TC Eugene for the SDM to be able to see it.

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: FNMOC web site had labels for some of its products  
for Fernanda around 1700 UTC indicating a central pressure of 996 mb. This  
did not come from the NHC b-deck. Later products have a correct pressure of  
1005 mb from the b-deck.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170712/fa634f83/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list